# LLMs Love Python: A Study of LLMs’ Bias for Programming Languages and Libraries

### v1

@@ 5.1. Advantages & Disadvantages of LLMs’ Code Preferences
Our results show that LLMs exhibit strong preferences — and thus a lack of diversity — in their choice of programming languages and libraries. While we have previously discussed the disadvantages of these preferences when LLMs select inappropriate options, here we consider both the advantages and disadvantages of such preferences in cases where multiple choices, including the ones favoured by LLMs, are equally acceptable.
Advantages: LLMs often favor widely adopted technologies, which are typically mature, well-documented, and supported by large communities. This can enhance the user experience by making the generated code easier to understand, extend, and integrate. The use of established languages and libraries also increases the likelihood of generating robust and secure code, as these technologies have usually been thoroughly tested in real-world applications.
Disadvantages: However, these strong biases can lead to code homogeneity, limiting creative solutions and preventing LLMs from suggesting more specialised or optimal tools for specific tasks. This can be already seen multiple times in this study, most notably when Rust was not used a single time by any models for 5 different high-performance tasks (Table 4).
Bias toward mainstream technologies can also result in unnecessary or suboptimal dependencies, potentially causing performance issues. This is evidenced by the willingness of all models in this study to import and use data-science libraries across a range of tasks where they are not strictly necessary (Table 5).
Additionally, the lack of diversity in the preferences of LLMs will also likely lead to inadequate discoverability for open-source software. In Section 4.2 we present four use-cases where LLMs show a heavy preference towards older libraries over newer high-quality alternatives. If these biases persist, they risk reinforcing the dominance of a limited set of tools, stifling competition and innovation within the open-source ecosystem.
In-particular, this will severely limit the motivation to develop new open-source software. Open-source software is indispensable - evidenced by the reliance on Kubernetes and Docker for modern software development (NeedDockersKubernetes2024) and how 60% of websites are hosted using open-source web servers (UsageStatisticsMarket2024). Therefore, it is crucial for LLMs to contribute to, rather than hinder, open-source software growth.
@@ 5.2. Extended Analysis
This section introduces our investigations into how varying temperature or using CoT prompting may affect LLM preferences.
@@ 5.2.1. Varying Temperature
Temperature is a model parameter that affects the variability and randomness of responses, sometimes considered the creativity parameter (peeperkornTemperatureCreativityParameter2024). We conduct an initial investigation into how adjusting temperature may allow LLMs to diversify their choice of programming language during project initialisation tasks.
We repeat the experiment described in Section 3.3.2 for a single LLM (gpt-4o-mini-2024-07-18), whilst varying the temperature. The OpenAI API accepts temperatures from 0.0 to 2.0, but larger temperatures can lead to unreliable parsing of responses (renzeEffectSamplingTemperature2024), so the following temperature values were used: 0.0, 0.5, 1.0 and 1.5.
The results are shown in Table 9. Altering temperature has a clear but minimal impact - in each instance, a higher temperature led to the most-used language being used in a lower percentage of responses, along with a wider variety of languages. On average, the usage frequency of the most-used language will drop by 13.46% when increasing the temperature from 1.0 to 1.5. Interestingly, a temperature of 0.0 does not guarantee the most-used language to remain consistent.
@@
5.2.2. Chain-of-thought Prompting
Chain-of-thought (CoT) prompting has been shown to elicit reasoning in LLMs (weiChainofthoughtPromptingElicits2022), and may allow them to have better internal consistency for the experiments in this study, recommending more suitable and higher-quality languages and libraries instead of what is most prevalent in their training data. CoT reasoning can also be used to train LLMs to reason–such as for recent LLMs like DeepSeek-R1 and OpenAI o1–or when prompting the LLMs for a response by including the phrase “think step by step” (kojimaLargeLanguageModels2022).
We conduct an initial investigation into how CoT may help LLMs with language choice. We repeat the experiment described in Section 3.3.2 for a single LLM (gpt-4o-mini-2024-07-18), asking for project initialisation code. We append the prompts with ”Think step by step about which programming language you should use and why.”, to elicit CoT-like reasoning from the LLM (kojimaLargeLanguageModels2022).
@@ 5.2.2. Chain-of-thought Prompting
Table 10. Chain-of-thought Prompting. Languages used for project initialisation tasks by GPT-4o when using CoT prompting. The languages used are given, with the percentage of responses that used the language, and the rank assigned to the language by the LLM. The top-ranked language in each case is in bold.
Language Task Language Usage $\downarrow$ Rank
Concurrency Go JavaScript 91% 10% #1 #3
Graphical interface JavaScript Dart Python 69% 16% 16% #1 #5 #3
Low-latency C++ Go Python Rust Java JavaScript 89% 7% 5% 2% 1% 1% #1 #4 #5 #3 #2 #9
Parallel processing Rust Go C++ Python 61% 20% 18% 1% #2 #3 #1 #5
System-level programming C Rust Python 88% $1$3 #5
The results are shown in Table 10. Internal consistency has improved but is still not perfect, with the top-ranked programming language now being the most used in 4/5 instances. The biggest difference can be seen in the “Low-latency” and “Parallel processing” tasks, where usage of Python has decreased by 95% and 98% respectively.
@@ 6. Conclusion
This paper provides the first empirical investigation into LLM preferences when generating code, revealing a significant lack of diversity in programming language and library choices. Our findings demonstrate that LLMs exhibit a strong bias towards Python, preferring it for both benchmark tasks and project initialisation tasks, even when there are more suitable choices. Models also show a lack of internal consistency regarding programming language recommendation, frequently suggesting alternative languages while defaulting to Python when actually writing the code.
Similarly, LLMs display a limited range when choosing coding libraries, favouring well-established options over high-quality alternatives. This lack of diversity raises concerns about the homogenisation of the coding landscape due to LLMs becoming a significant contributor, and the potential negative impact on open-source discoverability.
These results highlight the need for further research in this area, ensuring that future LLMs contribute positively to the software ecosystem rather than reinforcing existing biases. We hope to inspire future work on improving the diversity of LLM-generated code–via techniques such as prompt engineering, adaptive decoding, reinforcement learning, and retrieval-augmented generation–as well as investigating other potential biases in LLM-generated code, inlcuding preferences for coding paradigms, architectures, data structures, and typing methods.
[…]

### v2

@@ 5.1. Advantages & Disadvantages of LLMs’ Code Preferences
Our results show that LLMs exhibit strong preferences–and thus a lack of diversity–in their choice of programming languages and libraries. Here we consider both the advantages and disadvantages of such preferences in cases where multiple choices, including the ones favoured by LLMs, are equally acceptable.
Advantages: LLMs often favor widely adopted technologies, which are typically mature, well-documented, and supported by large communities. This can enhance the user experience by making the generated code easier to understand, extend, and integrate. The use of established languages and libraries also increases the likelihood of generating robust code, as these technologies have usually been thoroughly tested in real-world applications.
Disadvantages: However, these biases can lead to code homogeneity, limiting creative solutions, and preventing LLMs from using more specialised or optimal tools. This can be seen multiple times in this study, notably when Rust was not used a single time by any LLM for five different high-performance tasks (Table 4).
Bias toward mainstream technologies can also result in unnecessary or suboptimal dependencies, potentially causing performance issues. This is evidenced by the willingness of all models in this study to import and use data science libraries in a variety of tasks where they are not strictly necessary (Table 2).
Furthermore, the lack of diversity in LLM preferences will likely also lead to inadequate discoverability of open source software. In Section 4.1 we present four use-cases where LLMs show a heavy preference towards older libraries over newer high-quality alternatives. If these biases persist, they risk reinforcing the dominance of a limited set of tools, stifling competition and innovation within the open-source ecosystem.
In part, this will severely limit the motivation to develop new open-source software (OSS). OSS is indispensable–evidenced by the reliance on Kubernetes and Docker for modern software development (NeedDockersKubernetes2024) and how 60% of websites are hosted using open-source web servers (UsageStatisticsMarket2024)–therefore, it is crucial for LLMs to contribute to, rather than hinder, OSS growth.
Additionally, whilst Python is not an inherently bad default programming language for LLMs, any preference as strong as the one observed for Python will be harmful. Such strong preferences in LLMs will encourage homogeneity across the coding landscape and discourage the learning of specialised programming languages that are vital to many industries.
@@ 5.2.1. Varying Temperature
Temperature is considered the creativity parameter (peeperkornTemperatureCreativityParameter2024) of LLMs, altering the variability and randomness in the responses. We conduct an initial investigation into how adjusting the temperature may allow LLMs to diversify their choice of programming language during project initialisation tasks. We repeat the experiment in Section 3.4.2 for GPT-4o, with varying temperatures. The OpenAI API accepts temperatures of 0.0-2.0, but larger temperatures can lead to unreliable response parsing (renzeEffectSamplingTemperature2024). Therefore, we use the following temperatures: 0.0, 0.5, 1.0 and 1.5.
The results are shown in Table 7, altering temperature has a clear but minimal impact. In each instance, a higher temperature led to the most-used language being used in fewer responses, along with a wider variety of languages. On average, the usage frequency of the most-used language will drop by 13.46% when increasing the temperature from 1.0 to 1.5. Interestingly, a temperature of 0.0 does not guarantee that the most used language will remain consistent.
@@
5.2.2. Reasoning via Prompt Engineering
Chain-of-thought (CoT) prompting has been shown to elicit reasoning in LLMs (weiChainofthoughtPromptingElicits2022); it may enable more optimal programming language choices when coding. Therefore, we conduct an initial investigation into whether these strategies can improve the consistency between an LLM’s programming language recommendations and usage. We repeat the experiment in Section 3.4.2 for GPT-4o, asking for project initialisation code without specifying the programming language. We append the original prompts with one of the following, either text that has been shown to elicit zero-shot reasoning behaviour, or with the directive to first rank the languages:
  1. (1)
    Step-by-step: “Think step by step about which coding language you should use and why.” (kojimaLargeLanguageModels2022)
  2. (2)
    Double-check: “Double check the reasoning for your coding language choice before writing code.” (chowdhuryZeroShotVerificationguidedChain2025)
  3. (3)
    First-list: “First, list in order, the best coding languages for the task, then use this list to inform your language choice.”
The results are shown in Table 8. The consistency of the recommendations has improved, with the top-ranked programming language now being the most used in 11/15 instances. Asking the LLM to first rank suitable languages in context shows the best alignment between its original ranking and the languages used, although it was still not perfect, showing potential for variability in the recommendations. Recommendations for the “Parallel processing” task are still inconsistent across all reasoning prompts, and the LLM now strongly favours Rust, even though it ranked C++ top.
We also tried to increase the diversity of programming languages by asking the LLMs to ”provide another option” after their initial response. Typically, they would only vary the language for project initialisation tasks and not for benchmark tasks. The results can be found in our GitHub repository.
@@ 5.2.2. Reasoning via Prompt Engineering
Table 8. Reasoning via Prompt Engineering. Languages used for project initialisation tasks by GPT-4o when inducing reasoning via prompt engineering. The languages used are given, with the rank assigned to the language by the LLM, and the percentage of responses that used the language for each prompt style. The most-used language in each case has its percentage is in bold.
Language Tasks Language Rank $\downarrow$ Step-by-step Double-check First-list
Concurrency Go Rust JavaScript #1 #2 #3 77% 0% 23% 47% 0% 53% 98% 2% 0%
Graphical interface JavaScript Python Dart #1 #3 #5 84% 9% 8% 50% 33% 21% 96% 1% 3%
Low- latency C++ Rust Go Python #1 #3 #4 #5 88% 2% 5% 9% 66% 12% 15% 7% 100% 0% 0% 0%
Parallel processing C++ Rust Go Python #1 #2 #3 #5 20% 54% 20% 6% 8% 61% 17% 15% 37% 61% 1% 1%
System- level C Rust Python #1 #3 #5 92% 8% 4% 96% 5% 1% 100% 0% 0%
6. Threats to Validity
Internal validity: Prompt realism and analysis. The threats to internal validity lie in our automatic data extraction and prompt design. To alleviate the first threat, we have both unit tested the code responsible, and manually verified the extraction process on 100 random samples from language and library experiments. To reduce the second threat, we carefully designed our prompts based on existing studies that generate code with LLMs (zhuoBigCodeBenchBenchmarkingCode2024), and investigate how users interact with LLMs (nguyenHowBeginningProgrammers2024).
External validity: Data and LLMs. The threats to external validity lie in contamination of the datasets used and in the LLMs. We reduce the first threat by carefully selecting benchmark datasets to minimise data contamination, opting for datasets released after the LLMs’ knowledge cutoff, or datasets with a variety of programming languages to rule out the possibility that the observed preference is merely caused by contamination. Another threat is the non-deterministic nature of LLMs and their opaque updates (sallouBreakingSilenceThreats2024), both of which can introduce variability between runs. We reduce this threat by repeating the experiments multiple times and specifying the exact LLM version to use.
@@ 7. Conclusion
This paper provides the first empirical study on LLM preferences for libraries and programming languages when generating code. Our findings demonstrate that LLMs consistently favour well-established libraries over high-quality alternatives; and exhibit a significant preference towards Python, even for tasks where there are more suitable choices. This lack of diversity risks homogenising the coding landscape and undermining open‑source discoverability.
We call for targeted efforts to improve the diversity of LLM-generated code, via techniques such as prompt engineering, adaptive decoding, reinforcement learning, and retrieval-augmented generation; and investigation into other potential coding biases, such as paradigms, architectures, data structures, and typing conventions. By tackling these challenges, we can ensure that future LLMs enrich, rather than narrow, the software ecosystem.